QUESTION 5

For the below ME alternatives , which machine should be selected based on the PW analysis. MARR=10%.



Machine AMachine BMachine C
First cost, $              15000            30000          10,423
Annual cost, $/year                 12,340               6,000            4,000
Salvage value, $                 4,000               5,000            1,000
Life, years 362


Answer the below questions :

C- PW for machine C =
Answer #1

SOLUTION :


r = MARR = 10% = 0.1

=> (1 + r) = 1.1


PV of costs of Machine A :


= 15000 + 12340/ 1.1 + 12340/1.1^2 + 12340/1.1^3 - 4000/1.1^3 

= 42682.49 ($) 


PV of costs of Machine B :


= 30000 + 6000 (1.1^6 - 1)/(0.1*1.1^6) - 5000/1.1^6

= 53309.19 ($) 


PV of costs of Machine C :


= 10423 + 4000/1.1 + 4000/1.1^2 - 1000/1.1^2

= 16538.70 ($)


Therefore, as per PV of costs , Machine C has lowest cost. So Machine C should be selected.

(ANSWE

answered by: Tulsiram Garg